OpenStudy (anonymous):

Integrate ((ax+b)/((ax^2+2bx+c)^(1/2))dx

OpenStudy (dumbcow):

use substitution u = ax^2 +2bx+c du = 2ax +2b = 2(ax+b) dx \[\rightarrow \frac{1}{2} \int\limits_{}^{}\frac{1}{\sqrt{u}}du\]